OpenStudy (rainbow_dash):

Find the sum of the given infinite geometric series. 1 + 2/3 + 4/9 + ...

OpenStudy (anonymous):

you are supposed to recognize this as \(1+r+r^2+r^3+...=\sum_{k=1}^{\infty}r^{k-1}=\frac{1}{1-r}\)

hartnn (hartnn):

can u find the common ratio between them ?

OpenStudy (rainbow_dash):

ok.

OpenStudy (rainbow_dash):

i never covered this before in my real math. and the guys on this site don't explain anything, so sorry if i ask the same question in a different format

OpenStudy (anonymous):

so as @hartnn said, your actual job is to recognize \(r\) then it is a simple computation like the last one

OpenStudy (anonymous):

all of these questions are about summing a "geometric series" geometric series means something of the form \[a+ar+ar^2+ar^3+...\] where you multiply one term by the common ratio \(r\) to get to the next term

OpenStudy (anonymous):

this is more succinctly written is "sigma notation" as \[a+ar+ar^2+ar^3+..=\sum_{k=1}^{\infty}ar^{k-1}\] or as \[\sum_{k=0}^{\infty}ar^k\]

OpenStudy (anonymous):

and if \(-1<r<1\) you can use \[\sum_{k=1}^{\infty}ar^{k-1}=\frac{a}{1-r}\]

OpenStudy (anonymous):

the reasoning behind it is not too hard to understand, but may be a bit hard to write here are you using a text?

OpenStudy (rainbow_dash):

No, this is online

OpenStudy (rainbow_dash):

but i've written down every equation and formula you've given me on here

OpenStudy (anonymous):

that figures but since you are on line if you want to know why the formula works, you can probably get lots of explanation by googling i would recommend looking here http://patrickjmt.com/

OpenStudy (anonymous):

on the other hand if you are content with the formula i wrote it above also feel free to post similar questions, no one will be annoyed, and if they are that is their problem

OpenStudy (rainbow_dash):

Thanks a ton dude :) I really appreciate the help. much better than my old math teacher could do

OpenStudy (anonymous):

yw